OpenStudy (cake):

What is the solution to the equation k - 4 3/4 = 8 1/4 k= 4 k= 4 1/2 k= 12 k= 13

OpenStudy (cake):

hi

OpenStudy (anonymous):

\[k- \frac{ 19 }{ 4 }=\frac{ 33 }{ 4 }\]

OpenStudy (cake):

it has to be one of the answers above

OpenStudy (anonymous):

\[k=13\]

OpenStudy (anonymous):

\[4k-19=33\]

OpenStudy (cake):

r u sure

OpenStudy (anonymous):

\[4k=52\]

OpenStudy (anonymous):

13

OpenStudy (cake):

it k 13

OpenStudy (anonymous):

\[k=\frac{ 52 }{ 4}=13\]

OpenStudy (anonymous):

are you convinced now

OpenStudy (anonymous):

i know that k=13